# Gravity as a topological gauge theory

### v3

@@ Abstract
We describe a topological gauge theory with underlying Cartan geometry $G/H$ and principal $H$-bundle. The topological action is constructed as a linear combination of the Euler and Pontrjagin numbers coming from entries of the curvature $\bar{\Omega}$ of a Cartan connection.
As an example we show that for $G=SO(4,1)$ (respectively $SO(3,2)$) for $\Lambda_{0}>0$ (respectively $\Lambda_{0}<0$), and $H=SO(3,1)$, a very particular linear combination yields the Holst + Euler and Pontrjagin of the curvature $R$ of the spin connection + Nieh-Yan + bare cosmological constant $\Lambda_{0}$ terms. Additionally, in this construction, the coupling constants of these different terms are inherently linked together.
The other example treated is the one of the Möbius group $G=SO(4,2)/\{\pm I\}$ and $H=CO(4,2)\ltimes\mathbb{R}^{*3,1}$. By considering a constraint on the "pair of frames" $(\alpha,\beta)$ we retrieve an action comprising all terms compatible with diffeomorphism invariance and local Lorentz invariance described in the previous example + kinetic term for a scalar field $+$ an interaction of that scalar field with the Torsion $T$, the scalar field being related to dilations. Finally, in both examples, we study the equations of motion associated to the total (Gauge + Matter) action. It is especially shown that one of the solutions in the Möbius case consists of Einstein’s equations modified by the Holst and Nieh-Yan terms with an additional source term for curvature depending on specific variations of spin density of matter and on torsion.
Keywords : Topological gravity, Cartan geometry, Characteristic classes, Conformal gravity
@@ 1 Introduction
There exist several approaches to describe gravity as a gauge theory. Two of them are MacDowell-Mansouri gravity and BF theory, see e.g. [1]. They can be mathematically described in terms of Cartan geometries $G/H$ of manifolds $\mathscr{M}$ where $G,H\subset G$ and $\mathfrak{g},\mathfrak{h}$ are respectively Lie groups and their corresponding Lie algebras. When the geometry is reductive, the associated $\mathfrak{g}$-valued Cartan connection $\varpi$ naturally splits into two parts $\varpi=\varpi_{\mathfrak{h}}\oplus\varpi_{\mathfrak{p}}$, one with values in $\mathfrak{h}$ and the other one in $\mathfrak{p}=\mathfrak{g}/\mathfrak{h}$. In these formalisms the $\mathfrak{h}$-valued part $\varpi_{\mathfrak{h}}$ corresponds to the spin connection $A$ while the $\mathfrak{p}$-valued part $\varpi_{\mathfrak{p}}=\sqrt{\dfrac{\Lambda}{3\epsilon}}\,\beta$ is proportional to the vielbein 1-form $\beta$ (tetrad if $dim(\mathscr{M})=4$).
For a resulting Lorentzian [^3] Riemaniann or Galilean spacetimes can also be described by this formalism. For more details see [1, p.10]. The kind of topological gauge theories developed in this article are also compatible with these choices of Lie groups. space-time the structure group is given by $H=SO(3,1)$ while the Lie group $G$ is chosen according to the value and sign of the cosmological constant $\Lambda$ as in [1, p.10]:
$\left\{\begin{array}[]{lcll}G=SO(4,1)&\text{ and }&\epsilon=+1,&\text{ for }\Lambda>0\\ G=SO(3,2)&\text{ and }&\epsilon=-1,&\text{ for }\Lambda<0.\end{array}\right.$ (3)
@@
$(c)~{}\varpi(\mathfrak{X})\neq 0\text{ for every nonzero vector field }\mathfrak{X}\in\text{Vect}(\mathscr{P}).$
@@
For $m=4$, the MacDowell-Mansouri action is given by:
$S_{MM}[\varpi]$ $=-\dfrac{3}{2G\Lambda}\int_{\mathscr{M}}\operatorname{tr}(\bar{\Omega}_{\mathfrak{h}}\wedge\star\bar{\Omega}_{\mathfrak{h}})=-\dfrac{3}{4G\Lambda}\int_{\mathscr{M}}\varepsilon_{abcd}\,\bar{\Omega}_{\mathfrak{h}}^{ab}\wedge\bar{\Omega}_{\mathfrak{h}}^{cd}\ .$ (4)
This action is $Ad(H)$-invariant, i.e. Lorentz invariant.
@@
where $\operatorname{Pf}(\Omega_{\mathfrak{so}})$ is the Pfaffian of the $\mathfrak{so}(m)$-valued curvature 2-form $\Omega_{\mathfrak{so}}$. This is a $SO(m)$-invariant polynomial.
@@
Up to an overall factor, one can recognise the MacDowell-Mansouri action (4) which is recast as the Euler number associated to the $\mathfrak{h}=\mathfrak{so}(3,1)$-valued part $\bar{\Omega}_{\mathfrak{h}}$ of the curvature and which is $SO(3,1)$-invariant. This provides a very clear mathematical interpretation to the MacDowell-Mansouri action:
$S_{MM}[\varpi]=-\dfrac{24\pi^{2}}{G\Lambda}\,\mathdutchcal{E}$ (6)
Another characteristic class which will be of interest in the remainder of the article is the Pontrjagin characteristic class. As described in [4], one has:
@@
Let $L$ be a $k$-dimensional real vector bundle over an $m$-dimensional manifold $\mathscr{M}$. $\Omega$ is the curvature associated to a connection $\omega$ on the principal $H$-bundle over $\mathscr{M}$. Introducing orthonormal frames allows to reduce the structure group to $O(k)$ and one has $\mathfrak{o}(k)=\mathfrak{so}(k)$ at the level of Lie algebras. The $\mathfrak{so}(k)$-valued curvature is written $\Omega_{\mathfrak{so}}$. The corresponding total Pontrjagin class is then defined as: $P(\Omega_{\mathfrak{so}})=\det\Bigl{(}\mathds{1}+\dfrac{\Omega_{\mathfrak{so}}}{2\pi}\Big{)}$.
@@
$\mathdutchcal{P}=\int_{\mathscr{M}}P(\mathscr{M})=\int_{\mathscr{M}}P_{1}(\mathscr{M})=\int_{\mathscr{M}}-\frac{1}{8\pi^{2}}\operatorname{Tr}(\Omega_{\mathfrak{so}}\wedge\Omega_{\mathfrak{so}})\ ,$ (7)
where $P_{1}$ is the $1$st Pontrjagin class.
As stated in [4, Chern-Weil theorem], a nice feature of characteristic classes is that up to an exact term, they do not depend on the choice of connection.
Section 2 will introduce and describe the topological gauge theories associated to the Lie groups $G=SO(4,1),SO(3,2)$ considered as "mutations" (according to [5]) of the group $G=ISO(3,1)$. All these three groups contain $H=SO(3,1)$ which will be the structure group of the principal bundle $\mathscr{P}$. Then, a topological action as a linear combination of the Euler and Pontrjagin numbers will be considered for these Lie groups $G$ and $H$. Results of Section 2 are easily applicable to other spacetimes whether Riemaniann or Galilean or entirely different Lie groups.
In particular, Section 3 will be devoted to the construction of a similar topological action for the case where $G=SO(4,2)/\{\pm I\}$ (the Möbius group) and $H=CO(4,2)\ltimes\mathbb{R}^{*3,1}$, with $CO(p,q)$ the conformal group of signature $(p,q)=(4,2)$ here.
@@
2 Characteristic polynomials for a Cartan connection with Lorentzian spacetimes
@@ 2.1 Lorentzian (dS and AdS) spacetimes as mutations of an
Let $G$, $H\subset G$ be Lie groups with corresponding Lie algebras $\mathfrak{g}=\mathfrak{h}\oplus\mathfrak{p}$ such that $G/H$ is reductive, that is that the splitting $\mathfrak{g}=\mathfrak{h}\oplus\mathfrak{p}$ is $Ad(H)$-invariant.
One such reductive model geometry is $G/H=ISO(3,1)/SO(3,1)$, with algebra $\mathfrak{g}=\overbrace{\mathfrak{so}(3,1)}^{\mathfrak{h}}\oplus\overbrace{\mathfrak{\mathbb{R}}^{3,1}}^{\mathfrak{p}}$ where $G=ISO(3,1)=SO(3,1)\ltimes\mathbb{R}^{3,1}$ and $H=SO(3,1)$ is defined by the relation
$\Lambda^{\text{T}}\eta\Lambda=\eta,\forall\Lambda\in SO(3,1)$ with $\eta$ the Minkowski metric:
@@
A group element $g\in G=ISO(3,1)$ is parametrized as ${\displaystyle g=\begin{pmatrix}\Lambda&\dfrac{1}{\ell}v\\ 0&1\end{pmatrix}}$ where $v\in\mathbb{R}^{3,1}$.
Linearization leads to:
$\varphi$ $=\begin{pmatrix}\varphi_{\mathfrak{h}}&\varphi_{\mathfrak{p}}\\ 0&0\end{pmatrix}\in\mathfrak{iso}(3,1)\text{ where }\varphi_{\mathfrak{h}}\in\mathfrak{so}(3,1)~{},~{}\varphi_{\mathfrak{h}}^{\text{T}}\eta+\eta\varphi_{\mathfrak{h}}=0~{},~{}\varphi_{\mathfrak{p}}=\dfrac{1}{\ell}\nu\in\mathbb{R}^{3,1}$
@@
In [5, p218], Sharpe introduces the notion of mutation of a Cartan geometry $G/H$ according to the
@@
Let $(\mathfrak{g},\mathfrak{h})$ and $(\mathfrak{g}^{\prime},\mathfrak{h})$ be two model geometries. A mutation map corresponds to an $Ad(H)$ module isomorphism $\mu:\mathfrak{g}\rightarrow\mathfrak{g}^{\prime}$ (i.e. $\mu(Ad_{h}(u))=Ad_{h}(\mu(u))\,$) satisfying:
$(i)~{}\mu_{|\mathfrak{h}}=\mathrm{id}_{\mathfrak{h}}$
@@
A particular mutation that will be relevant for us is the one where $(\mathfrak{g},\mathfrak{h})=\bigr{(}\mathfrak{iso}(3,1),\mathfrak{so}(3,1)\bigr{)}$ and where the mutation map $\mu:\mathfrak{g}\rightarrow\mathfrak{g}^{\prime}$ is given by  [^5] More general mutations can be considered, e.g. $\varphi_{\mathfrak{g}}=\begin{pmatrix}\varphi_{\mathfrak{h}}&\varphi_{\mathfrak{p}}\\ 0&0\end{pmatrix}\longmapsto\varphi_{\mathfrak{g}^{\prime}}=\mu(\varphi_{\mathfrak{g}})=\begin{pmatrix}\varphi_{\mathfrak{h}}&k\varphi_{\mathfrak{p}}\\ k^{\prime}\varphi_{\bar{\mathfrak{p}}}&0\end{pmatrix}\in\mathfrak{g}^{\prime}$. However, whenever $k$ and $k^{\prime}$ are constants, this other mutation doesn’t change the physics that we will describe later on.
$\varphi_{\mathfrak{g}}=\begin{pmatrix}\varphi_{\mathfrak{h}}&\varphi_{\mathfrak{p}}\\ 0&0\end{pmatrix}\longrightarrow\varphi_{\mathfrak{g}^{\prime}}=\mu(\varphi_{\mathfrak{g}})=\begin{pmatrix}\varphi_{\mathfrak{h}}&\varphi_{\mathfrak{p}}\\ k\varphi_{\bar{\mathfrak{p}}}&0\end{pmatrix}\in\mathfrak{g}^{\prime}$ (1)
with $\varphi_{\bar{\mathfrak{p}}}=\varphi_{\mathfrak{p}}^{\text{T}}\eta$ and where $k$ may a priori be considered as 00-form scalar field.
@@
$\left\{\begin{array}[]{lcll}k=-1\Leftrightarrow G=SO(4,1)&\text{ and }&\epsilon=+1,&\text{ for }\Lambda_{0}>0,\\ &&&\\ k=+1\Leftrightarrow G=SO(3,2)&\text{ and }&\epsilon=-1,&\text{ for }\Lambda_{0}<0.\end{array}\right.$ (5)
Throughout the paper $k$ will be considered as a real number, the case of a 00-form scalar field $k$ will be studied in a forthcoming paper.
On the other hand, the canonical decomposition of a symmetric algebra $\mathfrak{g}=\mathfrak{h}\oplus\mathfrak{m}$ (see [6, Prop.2.1, Chapter XI] for more examples) corresponds to the commutators:
@@
Therefore, one can see that the mutated Lie algebra $\mathfrak{g}^{\prime}=\mathfrak{h}\oplus\mathfrak{m}$ of (1) corresponds to a symmetric algebra for $k\neq 0$, as a particular case of a reductive geometry with $\mathfrak{m}=\mathfrak{g}^{\prime}/\mathfrak{h}\simeq\mathfrak{p}=\mathfrak{g}/\mathfrak{h}$. Moreover, the generators of $\mathfrak{h}=\mathfrak{so}(3,1)$, and elements of the basis of $\mathfrak{p}=\mathfrak{\mathbb{R}}^{3,1}$, $\mathfrak{p}^{*}=\mathfrak{\mathbb{R}}^{3,1*}$ and ${\bar{\mathfrak{p}}}=\bar{\mathbb{R}}^{3,1}$ obey the relations (the isomorphism $\mathrm{End}(\mathbb{R}^{n})\simeq\mathbb{R}^{n}\otimes{\mathbb{R}^{n*}}$ is used):
@@
with $C^{fe}_{ab,cd}=-C^{ef}_{ab,cd}$ which is easily computed.
@@
$\varphi$ $=\begin{pmatrix}\varphi_{\mathfrak{h}}&\varphi_{\mathfrak{p}}\\ k\varphi_{\bar{\mathfrak{p}}}&0\end{pmatrix}=\begin{pmatrix}\dfrac{1}{2}\varphi_{\mathfrak{h}}^{ab}J_{ab}&\varphi_{\mathfrak{p}}^{a}e_{a}\\ k\varphi_{p}^{a}\bar{e}_{a}&0\end{pmatrix}=\dfrac{1}{2}\varphi_{\mathfrak{h}}^{ab}\begin{pmatrix}J_{ab}&0\\ 0&0\end{pmatrix}\oplus\varphi_{\mathfrak{m}}^{a}\begin{pmatrix}0&e_{a}\\ k\bar{e}_{a}&0\end{pmatrix}$ (6)

### v5

# Gravity as a deformed topological gauge theory

@@ Abstract
We describe gauge theories which allow to retrieve a large class of gravitational theories, including, MacDowell-Mansouri gravity and its topological extension to Loop Quantum Gravity via the Pontrjagin characteristic class involving the Nieh-Yan term. Considering symmetric spaces parametrized by mutations allows to naturally obtain a bare cosmological constant which in particular cases gives rise to a positive effective cosmological constant while having an AdS spacetime.
Two examples are studied, Lorentzian geometry (including dS and AdS spacetimes) and Lorentz×$\times$×Weyl geometry. In the latter case, we prove the equations of motion exhibit a secondary source for curvature in addition to the usual energy-momentum tensor. This additional source is expressed in terms of the spin density of matter, torsion and their variations.
Finally, we show that the gauge + matter actions constructed from invariant polynomials are asymptotically topological if one assumes a vanishing bare cosmological constant together with gauge and matter fields having compact support.
Keywords: Topological gravity, Homogeneous spaces, Cartan geometry, Characteristic classes
@@ 1 Introduction
There exist several approaches to describe gravity as a gauge theory. Two of them are MacDowell-Mansouri gravity and BF theory, see e.g. [1]. They can be mathematically described in terms of Cartan geometries on a manifold $\mathscr{M}$ modeled on a homogeneous space $G/H$ where $G,H\subset G$ and $\mathfrak{g},\mathfrak{h}$ are respectively Lie groups and their corresponding Lie algebras. When the geometry is reductive, namely, $\mathfrak{g}=\mathfrak{h}\oplus\mathfrak{p}$, a Ad$(H)$-invariant decomposition, the associated $\mathfrak{g}$-valued Cartan connection $\varpi$ naturally splits into two parts $\varpi=\varpi_{\mathfrak{h}}\oplus\varpi_{\mathfrak{p}}$, one with values in $\mathfrak{h}$ and the other one in $\mathfrak{p}=\mathfrak{g}/\mathfrak{h}$. In these formalisms the $\mathfrak{h}$-valued part $\varpi_{\mathfrak{h}}$ corresponds to the spin connection $A$ while the $\mathfrak{p}$-valued part is arranged to be $\varpi_{\mathfrak{p}}=\sqrt{\dfrac{\Lambda_{0}}{3\epsilon}}\begin{pmatrix}0&\beta\\ -\epsilon\bar{\beta}&0\end{pmatrix}$ is proportional to the vielbein 1-form $\beta$ (tetrad if $dim(\mathscr{M})=4$) where $\bar{\beta}=\eta(\beta)$ with $\eta$ the Minkowski metric.
For a resulting Lorentzian [^3] Riemaniann or Galilean spacetimes can also be described by this formalism. For more details see [1, p.10]. The kind of topological gauge theories developed in this article are also compatible with these choices of Lie groups. space-time the structure group is given by $H=SO(3,1)$ while the Lie group $G$ is chosen according to the value and sign of the (bare) cosmological constant $\Lambda_{0}$ as in [1, p.10]:
$\left\{\begin{array}[]{lcll}G=SO(4,1)&\text{ and }&\epsilon=+1,&\text{ for }\Lambda_{0}>0\\ G=SO(3,2)&\text{ and }&\epsilon=-1,&\text{ for }\Lambda_{0}<0.\end{array}\right.$ (3)
@@
$(c)~{}\varpi(X)\neq 0\text{ for every nonzero vector field }X\in\text{Vect}(\mathscr{P}).$
In particular, condition (c) implies that tangent spaces to $\mathscr{M}$ are isomorphic to the quotient $\mathfrak{g}/\mathfrak{h}$ of dimension $m$.
@@
For $m=4$, the MacDowell-Mansouri action [1] is given by:
$S_{MM}[\varpi]$ $=-\dfrac{3}{2G\Lambda_{0}}\int_{\mathscr{M}}\operatorname{tr}(\bar{\Omega}_{\mathfrak{h}}\wedge\star\bar{\Omega}_{\mathfrak{h}})=-\dfrac{3}{4G\Lambda_{0}}\int_{\mathscr{M}}\varepsilon_{abcd}\,\bar{\Omega}_{\mathfrak{h}}^{ab}\wedge\bar{\Omega}_{\mathfrak{h}}^{cd}$ (4)
where $\bar{\Omega}_{\mathfrak{h}}=R-\dfrac{\Lambda_{0}}{3}\beta\wedge\bar{\beta}$. This action is $Ad(H)$-invariant, i.e. Lorentz invariant.
@@
where $\operatorname{Pf}(\Omega_{\mathfrak{so}})$ is the Pfaffian of the $\mathfrak{so}(m)$-valued curvature 2-form $\Omega_{\mathfrak{so}}$ originating from the $\mathfrak{so}(m)$-valued connection $1$-form $\omega_{\mathfrak{so}}$. This is an $SO(m)$-invariant polynomial.
@@
One can recognize the MacDowell-Mansouri action (4) is proportional to the Pfaffian of the $\mathfrak{h}=\mathfrak{so}(3,1)$-valued part $\bar{\Omega}_{\mathfrak{h}}$ of the curvature, which is $SO(3,1)$-invariant. This provides a very clear mathematical interpretation to the MacDowell-Mansouri action as the deformation of a Schwarz-type [^5] A more detailed account of topological gauge theories of Schwarz-type can be found in [5, 6, 7]. The Euler number also yields a topological field theory according to [8]. See also [9]. topological gauge theory:
$S_{MM}[\varpi]=-\dfrac{24\pi^{2}}{G\Lambda_{0}}\,\int_{\mathscr{M}}\operatorname{Pf}(\dfrac{\bar{\Omega}_{\mathfrak{h}}}{2\pi})$ (6)
where $\bar{\Omega}_{\mathfrak{h}}=R-\dfrac{\Lambda_{0}}{3}\beta\wedge\bar{\beta}$, and $-\dfrac{\Lambda_{0}}{3}\beta\wedge\bar{\beta}$ acts as the deformation of the $\bar{\Omega}_{\mathfrak{so}}=R$ curvature.
Considering Chern-Simons theory as an example of topological field theory of the Schwarz type, another characteristic class which will be of interest in this work is the Pontrjagin characteristic class. As described in [4], one has:
@@
Let $L$ be a $k$-dimensional real vector bundle over an $m$-dimensional manifold $\mathscr{M}$. $\Omega$ is the curvature associated to a connection $\omega$ on a principal $G$-bundle over $\mathscr{M}$. Introducing orthonormal frames allows to reduce the structure group to $O(k)$ and one has $\mathfrak{o}(k)=\mathfrak{so}(k)$ at the level of Lie algebras. The $\mathfrak{so}(k)$-valued curvature is written $\Omega_{\mathfrak{so}}$. The corresponding total Pontrjagin class is then defined as: $P(\Omega_{\mathfrak{so}})=\det\Bigl{(}\mathds{1}+\dfrac{\Omega_{\mathfrak{so}}}{2\pi}\Big{)}$.
@@
$\mathdutchcal{P}=\int_{\mathscr{M}}P(\Omega_{\mathfrak{so}})=\int_{\mathscr{M}}P_{1}(\Omega_{\mathfrak{so}})=\int_{\mathscr{M}}-\frac{1}{8\pi^{2}}\operatorname{Tr}(\Omega_{\mathfrak{so}}\wedge\Omega_{\mathfrak{so}})\ ,$ (7)
where $P_{1}$ is the $1^{\text{st}}$ Pontrjagin class.
As stated in [4, Chern-Weil theorem], a nice feature of characteristic classes is that up to an exact term, they do not depend on the choice of connection. This fact allows us to build actions that are invariant under the choice of connection.
Generalizing the observation made for (6), the gauge invariant actions studied in this work are given by the linear combination:
$S_{G}[\varpi]=\int_{\mathscr{M}}\big{(}rP(\bar{\Omega})+e\operatorname{Pf}(\dfrac{F}{2\pi})+y\det(\mathds{1}+\dfrac{\bar{\Omega}_{\mathfrak{h}}}{2\pi})\big{)}$ (8)
where the first term is the Pontryagin number of the manifold related to the $\mathfrak{g}$-valued Cartan curvature $\bar{\Omega}$ while the two other terms are only invariant polynomials corresponding to the Pfaffian $\operatorname{Pf}(\dfrac{F}{2\pi})$ and the determinant $\det(\mathds{1}+\dfrac{\bar{\Omega}_{\mathfrak{h}}}{2\pi})$. They are respectively evaluated on $F$ (the $\mathfrak{so}(m-1,1)$-valued part of the curvature $\bar{\Omega}_{\mathfrak{h}}$) and $\bar{\Omega}_{\mathfrak{h}}$ ($\mathfrak{h}$-valued) parts of the Cartan curvature.
If $\mathfrak{g}/\mathfrak{h}=\mathfrak{m}$ obeys the commutation relation $[\mathfrak{m},\mathfrak{m}]\subset\mathfrak{h}$ (symmetric Lie algebra), then $G/H$ is called a reductive homogeneous space [10]. In fact, this is the case for the construction given in [1] for the MacDowell-Mansouri action. We shall extend this idea, and we shall see that Cartan geometries modeled on reductive homogeneous spaces are intrinsically linked to the notion of cosmological constant in specific cases. In the examples given in this paper, reductive homogeneous spaces will be implemented at the infinitesimal level through symmetric Lie algebras. Using the concept of mutation, two symmetric Lie algebras where $H$ is either the Lorentz or Lorentz×$\times$×Weyl group will be given as resulting from a mutation procedure on which a Cartan geometry will be constructed along the line of [11].
The paper is organized as follows. Section 2 will introduce and describe the gauge theories associated to the Lie groups $G=SO(4,1)$ (respectively $SO(3,2)$) for $\Lambda_{0}>0$ (respectively $\Lambda_{0}<0$) considered as "mutations" (according to [11]) of the group $G=ISO(3,1)$. All these three groups contain $H=SO(3,1)$ which will be the structure group of the principal bundle $\mathscr{P}$. Then, the gauge invariant action (8) will be studied for these Lie groups $G$ and $H$. Subsenquently, a very particular linear combination yields the Holst + Euler and Pontrjagin of the curvature $R$ of the spin connection + Nieh-Yan + bare cosmological constant $\Lambda_{0}$ terms. Additionally, in this construction, the coupling constants of these different terms are inherently linked together. Results of Section 2 are adaptable to other spacetimes whether Riemannian or Galilean or for other Lie groups.
Section 3 will be devoted to the construction of a similar action for the case where $G=SO(4,2)/\{\pm I\}$ (the Möbius group) and $H=CO(3,1)=SO(3,1)\times(\mathbb{R}_{+}\setminus\{0\})$ the Lorentz×$\times$×Weyl group of signature $(3,1)$, so that the infinitesimal Klein geometry $(\mathfrak{g},\mathfrak{h})$ in the sense of [11] turns out to be a symmetric Lie algebra. Then, mutating the $(\mathfrak{g},\mathfrak{h})$ model allows to mimic what we did in the previous Lorentz case. Reducing the quotient $\mathfrak{g}/\mathfrak{h}$ to be of dimension $4$ by constraining the "pair of frames" $(\alpha,\beta)$ one can retrieve an action comprising all terms described in the previous example + a kinetic term for a scalar field (dilations).
In both examples, we study the equations of motion associated to the total action (gauge + matter). It is especially shown that, in the Möbius case, if one adds an interaction between dilations and torsion $T$, the equations of motion consist of Einstein’s equations modified by the Holst term with an additional source term for curvature depending on specific variations of spin density of matter and on torsion.
Lastly, in Section 4, by studying invariant polynomials for reductive homogeneous spaces, it will be shown that (upon restricting to connections with compact supports on $\mathscr{M}$) the mutated geometries can be considered as deformations such that the gauge invariant actions constructed in the previous sections are asymptotically invariant under the choice of this type of connections, and thus lead to topological theories.
@@
2 Invariant polynomials for a Cartan connection with Lorentzian spacetimes
@@ 2.1 Lorentzian (dS and AdS) spacetimes as mutations of an
Let $G$, $H\subset G$ be Lie groups with corresponding Lie algebras $\mathfrak{g}=\mathfrak{h}\oplus\mathfrak{p}$ such that $G/H$ is reductive [11], that is that the splitting $\mathfrak{g}=\mathfrak{h}\oplus\mathfrak{p}$ is $Ad(H)$-invariant.
One such reductive model geometry is $G/H=ISO(3,1)/SO(3,1)$, with algebra $\mathfrak{g}=\overbrace{\mathfrak{so}(3,1)}^{\mathfrak{h}}\oplus\overbrace{\mathfrak{\mathbb{R}}^{3,1}}^{\mathfrak{p}}$ where $G=ISO(3,1)=SO(3,1)\ltimes\mathbb{R}^{3,1}$ and $H=SO(3,1)$ is defined by the relation $\Lambda^{\text{T}}\eta\Lambda=~{}\eta,\forall\Lambda\in SO(3,1)$ with $\eta$ the Minkowski metric:
@@
A group element $g\in G=ISO(3,1)$ is parametrized as ${\displaystyle g=\begin{pmatrix}\Lambda&a\\ 0&1\end{pmatrix}}$ where $a\in\mathbb{R}^{3,1}$. Linearization leads to:
$\varphi$ $=\begin{pmatrix}\varphi_{\mathfrak{h}}&\varphi_{\mathfrak{p}}\\ 0&0\end{pmatrix}\in\mathfrak{iso}(3,1)\text{ where }\varphi_{\mathfrak{h}}\in\mathfrak{so}(3,1)~{},~{}\varphi_{\mathfrak{h}}^{\text{T}}\eta+\eta\varphi_{\mathfrak{h}}=0,~{}\text{and}~{}\varphi_{\mathfrak{p}}\in\mathbb{R}^{3,1}$
@@
In [11, see p.218], Sharpe introduces the notion of mutation of a Cartan geometry $G/H$ according to the
@@
Let $(\mathfrak{g},\mathfrak{h})$ and $(\mathfrak{g}^{\prime},\mathfrak{h})$ be two model geometries. A mutation map corresponds to an $Ad(H)$ module isomorphism $\mu:\mathfrak{g}\rightarrow\mathfrak{g}^{\prime}$ (i.e. $\mu(Ad_{h}(u))=Ad_{h}(\mu(u))\,\forall u\in\mathfrak{g}$) satisfying:
$(i)$ $\ \mu_{|\mathfrak{h}}=\mathrm{id}_{\mathfrak{h}}$
@@
In the case where the model geometry is $(\mathfrak{g},\mathfrak{h})=\bigr{(}\mathfrak{iso}(3,1),\mathfrak{so}(3,1)\bigr{)}$, a mutation that will be particularly relevant for us is given by the mutation map [^6] This mutation map generalizes the examples given in [11, Example 6.2, p.218]. $\mu:\mathfrak{g}\rightarrow\mathfrak{g}^{\prime}$:
$\varphi_{\mathfrak{g}}=\begin{pmatrix}\varphi_{\mathfrak{h}}&\varphi_{\mathfrak{p}}\\ 0&0\end{pmatrix}\mapsto\varphi_{\mathfrak{g}^{\prime}}=\mu(\varphi_{\mathfrak{g}})=\begin{pmatrix}\varphi_{\mathfrak{h}}&k_{1}\varphi_{\mathfrak{p}}\\ k_{2}\bar{\varphi}_{\mathfrak{p}}&0\end{pmatrix}\in\mathfrak{g}^{\prime}$ (1)
with $\bar{\varphi}_{\mathfrak{p}}=\varphi_{\mathfrak{p}}^{\text{T}}\eta$ and where $k_{1},k_{2}$ may a priori be considered as 00-form scalar fields. It is worthwhile to notice that this mutation transforms the geometry modeled on $G/H$ given earlier to the Lorentzian Cartan geometry modeled on $G^{\prime}/H$ with elements $g^{\prime}\in G^{\prime}$ obeying the group relation $g^{\prime T}Ng^{\prime}=N$, where $N$ is the following metric:
$N=\begin{pmatrix}\eta&0\\ 0&-k_{1}/k_{2}\end{pmatrix}$
@@
$\left\{\begin{array}[]{lcll}\dfrac{k_{1}}{k_{2}}=-1\Leftrightarrow G=SO(4,1)&\text{ and }&\epsilon=+1,&\text{ for }\Lambda_{0}>0,\\ &&&\\ \dfrac{k_{1}}{k_{2}}=+1\Leftrightarrow G=SO(3,2)&\text{ and }&\epsilon=-1,&\text{ for }\Lambda_{0}<0.\end{array}\right.$ (5)
Throughout the paper $k_{1}$ and $k_{2}$ will be considered as real numbers. It is however of physical interest to consider the mutation parameters as scalar fields, thus dealing with point-wisely dependent mutations; this will be studied in a forthcoming paper. Other papers considering dynamical versions of physical constants in the context of gravity include for instance [12, 13].
On the other hand, the canonical decomposition of a symmetric Lie algebra $\mathfrak{g}=\mathfrak{h}\oplus\mathfrak{m}$ (see [14, Prop.2.1, Chapter XI] for more examples) corresponds to the commutators:
@@
Therefore, one can see that the mutated Lie algebra $\mathfrak{g}^{\prime}=\mathfrak{h}\oplus\mathfrak{m}$ of (1) corresponds to a symmetric Lie algebra for $k\neq 0$, as a particular case of a reductive geometry with $\mathfrak{m}=\mathfrak{g}^{\prime}/\mathfrak{h}\simeq\mathfrak{p}=\mathfrak{g}/\mathfrak{h}$. Moreover, the generators of $\mathfrak{h}=\mathfrak{so}(3,1)$, and elements of the basis of $\mathfrak{p}=\mathfrak{\mathbb{R}}^{3,1}$, $\mathfrak{p}^{*}=\mathfrak{\mathbb{R}}^{3,1*}$ and ${\bar{\mathfrak{p}}}=\bar{\mathbb{R}}^{3,1}$ obey the relations (the isomorphism $\mathrm{End}(\mathbb{R}^{n})\simeq\mathbb{R}^{n}\otimes{\mathbb{R}^{n*}}$ is used):
@@
where the structure constants $C^{fe}_{ab,cd}=-C^{ef}_{ab,cd}$ can be easily computed.
@@
$\varphi$ $=\begin{pmatrix}\varphi_{\mathfrak{h}}&k_{1}\varphi_{\mathfrak{p}}\\ k_{2}\bar{\varphi}_{\mathfrak{p}}&0\end{pmatrix}=\begin{pmatrix}\dfrac{1}{2}\varphi_{\mathfrak{h}}^{ab}J_{ab}&k_{1}\varphi_{\mathfrak{p}}^{a}e_{a}\\ k_{2}\varphi_{\mathfrak{p}}^{a}\bar{e}_{a}&0\end{pmatrix}=\dfrac{1}{2}\varphi_{\mathfrak{h}}^{ab}\begin{pmatrix}J_{ab}&0\\ 0&0\end{pmatrix}\oplus\varphi_{\mathfrak{m}}^{a}\begin{pmatrix}0&k_{1}e_{a}\\ k_{2}\bar{e}_{a}&0\end{pmatrix}$ (6)